This is a basic question but I can seem to explain it. The Quadratic Equation term "b^2 -4ab" why do we change it to "b^2+ 4ab" when the "b" value is negative? Thanks annmrj@aol.com
EX: b^2-4ab (x^2-6x-27=0) = 6^2+4(1)(-27)=144

Answer by AnlytcPhil(1806)

This is a basic question but I can seem to explain it.
The Quadratic Equation term "b^2 -4ab" why do we change
it to "b^2+ 4ab" when the "b" value is negative?
----------------------------------------------------
We don't. First of all it's not b²-4ab, it's b²-4ac.
Also your example:
b^2-4ab (x^2-6x-27=0) = 6^2+4(1)(-27)=144
is incorrect in the next to last step, although 144 is correct.
It should be
b^2-4ab (x^2-6x-27=0) = 6^2-4(1)(-27)=144
What you have there, namely,
6^2+4(1)(-27)=
equals to -72, not 144.
Suppose we have
2x² + 5x - 3
then a=2, b=5, and c=-3
So
b² - 4ac
becomes
(5)² - 4(2)(-3)
or
25 - (-24)
and then the " - " sign before the parentheses and the " - "
before the 24 inside the parentheses becomes a " + ", and we
have
25 + 24
or
49
Edwin